OpenStudy (anonymous):

solve for -115 -10x = 65 +2x

OpenStudy (davidusa):

x=-15

OpenStudy (anonymous):

x=-15

OpenStudy (anonymous):

you move all the x value on one side and number on another side as: -115-65=10x+2x then -180=12x x=-180/12=-15 make sence ^^